OpenStudy (jfruscianterhcp):

given triangle ABC whose sides are AB = 15 in. AC 25 in. and BC 30 in. . from a point D on side AB, a line DE is drawn to a point E on side AC such that angle ADE is equal to angle ABC. If the perimeter of triangle ADE is 28 in. in find the lengths of line segments and CE?

OpenStudy (rane):

do u have the diagram?

OpenStudy (jfruscianterhcp):



ganeshie8 (ganeshie8):

you knw the perimeter of ADE find the perimeter of ABC also and take ratio of both, you wud get scale factor

OpenStudy (jfruscianterhcp):

how

ganeshie8 (ganeshie8):

how for what question ?

OpenStudy (jfruscianterhcp):

how to answer the problem show solution pls and explain

ganeshie8 (ganeshie8):

did u take time to read my reply above ?

ganeshie8 (ganeshie8):

find the perimeter of ABC first

OpenStudy (jfruscianterhcp):

well for its perimeter just add the three right?

OpenStudy (jfruscianterhcp):

so basically its 70

OpenStudy (jfruscianterhcp):

then?

ganeshie8 (ganeshie8):

Correct !

ganeshie8 (ganeshie8):

scale factor = 70/28 = 5/2

ganeshie8 (ganeshie8):

did u get, this far ?

OpenStudy (jfruscianterhcp):

yeah thats exactly what i did

OpenStudy (jfruscianterhcp):

wer my question

OpenStudy (jfruscianterhcp):

but it doesnt answer my question

ganeshie8 (ganeshie8):

well you know the scale factor, you have sides of one triangle. you can find the sides of other triangle, by multiplying the scale factor

OpenStudy (jfruscianterhcp):

how? I cant understand

ganeshie8 (ganeshie8):

btw, scale factor = 28/70 = 2/5

ganeshie8 (ganeshie8):